I buy a tank of gas for $41.56. That price is the actual cost of the gas plus
6.25% sales tax. How much sales tax did I pay? How much was the actual price of the gas?

Answer by htmentor(1343)

The price paid is the cost of the gas plus the tax
If c = the cost of the gas then the tax paid = 0.0625*c
So the total price = c + 0.0625c = 41.56
c(1+0.0625) = 41.56
c = 41.56/1.0625
c = 39.16
So the actual price of the gas is $39.12
The tax is 41.56 - 39.12 = $2.44